OpenStudy (anonymous):

A video game decreased in price from $50 to $45. What was the approximate percent decrease in the price? 0.1% 1% 5% 10%

Parth (parthkohli):

\[\Large \color{purple}{\rightarrow {Decrease \over Original} * 100 }\]

OpenStudy (anonymous):

5%

Parth (parthkohli):

No, \[\Large \color{purple}{\rightarrow {5 \over 50} * 100 }\]

OpenStudy (callisto):

% decrease= (50-45)/50 x 100% = 5/50 x 100% =1/10 x 100% =?

Parth (parthkohli):

\[\Large \color{purple}{\rightarrow {1 \over 10} * 100 }\]

OpenStudy (anonymous):

1%

Parth (parthkohli):

Na, \[\Large \color{purple}{\rightarrow {1 \over 1} * 10 }\]

OpenStudy (callisto):

.... 100% /10 = (100/10)% =?

OpenStudy (anonymous):

10

Parth (parthkohli):

Yes!